# Voltage dependent first-principles barriers to Li transport within Li-ion battery Solid-Electrolyte-Interphases

### v1

@@ 3.1.2 LiO
We predict the SEI and anode-SEI interfacial barrier for Li transport through a LiO SEI as a function of voltage in Fig. 4. In many ways, the barriers here resemble what we have already seen for the LiF SEI. The SEI barrier dominates the response of the system to electronic voltage, absorbing 93% of the voltage drop on the system. This does represent a slight increase in the barrier distributed at charge traps at the Li/LiO interface, rising to 7%. This rise can be interpreted as the density of charge traps slightly increasing in comparison to LiF, (the LiO interfacial structure shown in Fig. 4a even looks messier, a good indication that the system is being disrupted from its normal equilibrium, which often leads to charge traps).
@@
An important difference, however, is the anode potential at which the SEI barrier switches from encourage Li transport from the electrolyte to the anode to encouraging this transfer. For the LiO SEI, this threshold potential is 2.53 V (Li/Li)), nearly 0.6 V higher than what we saw for LiF. This implies that a Li battery with a LiO SEI will encourage Li transport toward the anode at higher anode potentials (and thus, by Eq. 6, lower open circuit voltages) than a battery with a LiF SEI. Depending on the application of interest, this could potentially be used to increase the energy density of a charged battery, or conversely to lower the voltage requirements for full charging.
3.1.3 LiCO
@@
We predict the SEI and anode-SEI interfacial barrier for Li transport through a LiCO SEI as a function of voltage in Fig. 5. For the LiCO SEI, the threshold voltage of switching from a negative barrier encouraging Li ion transport toward the anode to a positive barrier discouraging transport is 1.28 V (Li/Li)), nearly 0.6 V lower than what we saw for LiF. The nature of the surface states remain largely the same though, with the bulk of the barrier (94%) taking place across the bulk SEI. Only (6%) of the barrier is captured at surface states.
@@ 3.1.4 Grain boundaries
We apply QCA and predict the SEI barrier as a function of electronic voltage, as shown if Fig. 6d. We find that the proportion of the total potential barrier taken up by interfacial states (3%) versus the proportion that is distributed throughout the SEI (97%) is almost identical to the split in the interfacial barrier for a pure crystalline LiF SEI. Similarly, the voltage threshold at which the potential barrier switches from helping Li ion progress toward the anode to hindering progress is almost identical at 1.87 V (Li/Li)). This demonstrates that the surface state properties which determine the potential barrier within the SEI are largely independent of any grain boundaries that may (and in fact likely do) exist between two chemistries that are the same.
Finally, we examine an SEI composed of differing chemistries: a LiF-LiO grain boundary, with the atomic structure shown in Fig. 7a-b. We find that the proportion of the total barrier taken up by the interfacial charge trap states (7% as shown in Fig. 7c) matches that of LiO relatively well. The SEI barrier similarly composes 93% of the total barrier. This indicates that in a grain boundary with multiple chemical compositions, the level of interfacial charge trapping will be largely determined by the composition with the highest interfacial trapping. The voltage threshold of crossover between encouraging Li ion transport from the electrolyte toward the anode, however, is 0.50 V (Li/Li)). The SEI barrier therefore favors Li transport toward the anode (charging) near these grain boundaries only at significantly lower anode potentials than either LiF (1.93 V Li/Li) or LiO (2.53 V Li/Li) single composition SEIs. At lower potentials, this translates to larger SEI potential drops favoring Li transport toward the anode. For instance, at a voltage of 0.2 V (Li/Li), our calculations lead us to predict a SEI potential drop of -1.51 V in a LiF SEI, -2.15 V in LiO, and only -0.28 V in the modeled LiF-LiO grain boundary. This indicates that the specific LiF-LiO grain boundary modeled makes migration of the anode significantly harder, with larger open circuit voltages required to achieve similar levels of SEI assistance to Li-ion transport.
@@ 3.2 Li diffusion barriers
We simulate the knock-off mechanism for interstitial movement in a LiF SEI, as depicted in Fig. 8a. We then change the potential of both the starting and transition state with QCA to find the potential dependent barrier, as shown in Fig. 8b. This approach differs from previous work that examines Li interstitial barriers that only looks at neutral bulk conditions and thus ignores the contributions of voltage such as Zheng et al., who find a barrier of 0.25 eV for this reaction 39. Using standard DFT NEB calculations, we predict a 0.12 eV barrier. This is already significantly different and may be explained by our inclusion of a Li-metal anode in addition to the LiF. Further, the QCA derived potential of both the transition state and the starting state of the reaction are not kept constant during typical NEB calculations. This means we need to adjust the potential of both images to match before subtracting the energies to derive the QCA barriers, as described in Sec. 2.3.
@@
where is the anode electrochemical potential of the system with respect to the Li/Li standard electrode. This voltage dependence is fairly unimpressive! However, it largely tracks with the local distribution of the SEI barrier as this reaction barrier is for a Li ion traveling a much smaller distance than the entire SEI and thus will face only a fraction of the total SEI barrier. As a concrete example, at a voltage of 0.2 V (Li/Li), we expect to see a SEI barrier of -1.51 eV across the entire 200 Å LiF SEI favoring Li ion transport to the anode. The interstitial Li ion in the reaction shown in Fig. 8a is traveling 0.5 Å in the direction perpendicular to the interface, and we can thus expect that the portion of the SEI barrier it encounters over this period would proportionally be -0.0038 eV. When we predict the diffusion barrier for this system using Eq. 9, we predict that the diffusion barrier will be lowered by -0.0035 eV from its “neutral” value at the flatband potential, nearly an exact match.
@@
We can thus say that the impact of altering the voltage of the electrode SEI will be minimal for any given Li diffusion barrier over a few Å. This means that the vast majority of diffusion barriers from the existing DFT literature can be used with relatively minimal modification. There will, however, be a small alteration of the diffusion barrier corresponding to the magnitude of the SEI barrier at a given voltage divided by the fraction of the SEI length that the reaction takes place (assuming the charge carrier concentration is similar to what we model here, resulting in linear potential distributions). Furthermore, while the voltage based reduction of the barrier for any given Li movement across one site is minimal, this reaction needs to be repeated hundreds of times to cross the entire SEI. This means that the total amount of energy needed for a Li ion to move across the entire SEI can be significantly reduced (or increased) by changing the electronic voltage.
Overall, these results suggest that the specific chemistry of a given anode/SEI as well as their interface plays the primary role of determining how difficult Li ion transport will be. Altering the voltage on the material, however, can have impacts on the total energy needed, and represents a useful tool for tuning a given devices performance.
@@ 4 Discussion
Our simulations are necessarily imperfect representations of real battery devices. While QCA extends the length scale examined perpendicular to the surface, it does not extend the lateral area explored, which is treated using typical DFT periodic boundary conditions. This means that simulations of individual SEIs such as LiF and LiO inherently assume that only these perfectly crystalline chemistries exist in the SEI. In reality, the surface is known to likely be a mixture of different SEI chemistries interacting with each other. While our simulations of grain boundaries begin to address this concern, they are necessarily still periodic and are thus assuming a repeating pattern of grain boundaries across the surface, often with quite high surface densities. Our work will serve as a stepping stone, however, for future multiscale research aiming to address the larger lateral surface.
@@
Another aspect that limits the potential real world accuracy of our calculations is the representation of the electrolyte. For ease of demonstrating QCA within these systems, we only simulated rather shallow monolayers of DOL electrolyte molecules along with the inorganice SEI, but organic SEI components based on reactions with the electrolyte are often observed as well 40. This will be addressed with future work including a realistic mix of electrolyte molecules and a full Guoy-Chapman-Stern analytic inclusion of the charge distribution. QCA represents a tractable method for including as much information about electrolytes as feasible into these calculations before resorting to relatively expensive AIMD.
@@ 5 Conclusion
Our QCA analysis reveals that the barrier for transporting Li through the SEI of a battery is dependent on the overall applied electronic voltage, and to large extent on the location of the Li ion in the SEI. This implies that controlling the chemical composition of the SEI and applied voltage may be a useful tool for modulating the rate of charging for a given battery systems. In general, this trend matches intuition as higher open circuit voltages lead to lower transport barriers, favoring easier charging. The threshold voltage at which Li ion movement is encouraged toward the anode versus away from the anode, however, is highly dependent on the specific SEI chemistry. We see that, at least in some cases, grain boundaries lower this potential threshold more than would be seen for the individual components. Furthermore, we find that this SEI potential drop does impact the individual Li diffuision barriers, although at small levels for movements of a few Å. These small differences can, however, lead to significant differences in the total energy required to move from the electrolyte to the anode.

### v2

@@ 3.1.2 Li2O
We predict the SEI and anode-SEI interfacial barrier for Li transport through a Li2O SEI as a function of voltage in Fig. 4. In many ways, the barriers here resemble what we have already seen for the LiF SEI. The SEI barrier dominates the response of the system to electronic voltage, absorbing 94% of the voltage drop on the system. This does represent a slight increase in the barrier distributed at charge traps at the Li/Li2O interface, rising to 6%. This rise can be interpreted as the density of charge traps slightly increasing in comparison to LiF.
@@
An important difference, however, is the anode potential at which the SEI barrier switches from encourage Li transport from the electrolyte to the anode to encouraging this transfer. For the Li2O SEI, this threshold potential is 2.53 V (Li/Li+)), nearly 0.6 V higher than what we saw for LiF. This implies that a Li battery with a Li2O SEI will encourage Li transport toward the anode at higher anode potentials (and thus, by Eq. 9, lower open circuit voltages) than a battery with a LiF SEI. Depending on the application of interest, this could potentially be used to increase the energy density of a charged battery, or conversely to lower the voltage requirements for full charging.
3.1.3 Li2CO3
@@
We predict the SEI and anode-SEI interfacial barrier for Li transport through a Li2CO3 SEI as a function of voltage in Fig. 5. For the Li2CO3 SEI, the threshold voltage of switching from a negative barrier encouraging Li ion transport toward the anode to a positive barrier discouraging transport is 1.28 V (Li/Li+)), nearly 0.6 V lower than what we saw for LiF. The nature of the surface states remain largely the same though, with the bulk of the barrier (94%) taking place across the bulk SEI. Only (6%) of the barrier is captured at surface states.
@@ 3.1.4 Grain boundaries
We apply QCA and predict the SEI barrier as a function of electronic voltage, as shown if Fig. 6d. We find that the proportion of the total potential barrier taken up by interfacial states (3%) versus the proportion that is distributed throughout the SEI (97%) is almost identical to the split in the interfacial barrier for a pure crystalline LiF SEI. Similarly, the voltage threshold at which the potential barrier switches from helping Li ion progress toward the anode to hindering progress is almost identical at 1.87 V (Li/Li+)). This demonstrates that the surface state properties which determine the potential barrier within the SEI are largely independent of any grain boundaries that may (and in fact likely do) exist between two chemistries that are the same.
Finally, we examine an SEI composed of differing chemistries: a LiF-Li2O grain boundary, with the atomic structure shown in Fig. 7a-b. We find that the proportion of the total barrier taken up by the interfacial charge trap states (7% as shown in Fig. 7c) matches that of Li2O relatively well. The SEI barrier similarly composes 93% of the total barrier. This indicates that in a grain boundary with multiple chemical compositions, the level of interfacial charge trapping will be largely determined by the composition with the highest interfacial trapping. The voltage threshold of crossover between encouraging Li ion transport from the electrolyte toward the anode, however, is 0.50 V (Li/Li+)). The SEI barrier therefore favors Li transport toward the anode (charging) near these grain boundaries only at significantly lower anode potentials than either LiF (1.93 V Li/Li+) or Li2O (2.53 V Li/Li+) single composition SEIs. At lower potentials, this translates to larger SEI potential drops favoring Li transport toward the anode. For instance, at a voltage of 0.2 V (Li/Li+), our calculations lead us to predict a SEI potential drop of -1.51 V in a LiF SEI, -2.15 V in Li2O, and only -0.28 V in the modeled LiF-Li2O grain boundary. This indicates that the specific LiF-Li2O grain boundary modeled makes migration of the anode significantly harder, with larger open circuit voltages required to achieve similar levels of SEI assistance to Li ion transport.
@@ 3.2 Li diffusion barriers
3.2.1 Li transport in LiF
We simulate the knock-off mechanism for interstitial movement in a LiF SEI, as depicted in Fig. 8a. We then change the potential of both the starting and transition state with QCA to find the potential dependent barrier, as shown in Fig. 8b. This approach differs from previous work that examines Li interstitial barriers that only looks at neutral bulk conditions and thus ignores the contributions of voltage such as Zheng et al., who find a barrier of 0.25 eV for this reaction 41. Using standard DFT NEB calculations, we predict a 0.12 eV barrier. This is already significantly different and may be explained by our inclusion of a Li-metal anode in addition to the LiF. Further, the QCA derived potential of both the transition state and the starting state of the reaction are not kept constant during typical NEB calculations. This means we need to adjust the potential of both images to match before subtracting the energies to derive the QCA barriers, as described in Sec. 2.3.
@@ 3.2.1 Li transport in LiF
where is the anode electrochemical potential of the system with respect to the Li/Li+ standard electrode. This voltage dependence is fairly unimpressive! However, it largely tracks with the local distribution of the SEI barrier as this reaction barrier is for a Li ion traveling a much smaller distance than the entire SEI and thus will face only a fraction of the total SEI barrier. As a concrete example, at a voltage of 0.2 V (Li/Li+), we expect to see a SEI barrier of -1.51 eV across the entire 200 Å LiF SEI favoring Li ion transport to the anode. The interstitial Li ion in the reaction shown in Fig. 8a is traveling 0.5 Å in the direction perpendicular to the interface, and we can thus expect that the portion of the SEI barrier it encounters over this period would proportionally be -0.0038 eV. When we predict the diffusion barrier for this system using Eq. 12, we predict that the diffusion barrier will be lowered by -0.0035 eV from its “neutral” value at the flatband potential, nearly an exact match.
@@
We can thus say that the impact of altering the voltage of the electrode SEI will be minimal for any given Li diffusion barrier over a few Å. There will, however, be a small alteration of the diffusion barrier corresponding to the magnitude of the SEI barrier at a given voltage divided by the fraction of the SEI length that the reaction takes place (assuming the charge carrier concentration is similar to what we model here, resulting in linear potential distributions). Furthermore, while the voltage based reduction of the barrier for any given Li movement across one site is minimal, this reaction needs to be repeated hundreds of times to cross the entire SEI. This means that the total amount of energy needed for a Li ion to move across the entire SEI can be significantly reduced (or increased) by changing the electronic voltage.
There is currently no experimental evidence to clearly distinguish between the reaction barriers shown due to the difficulty of individual atomic measurements. Future work using careful in-situ analysis with tools such as Kelvin Probe Force Microscopy and Quartz Crystal Microbalance measurements could be useful for helping narrow the range of theoretical uncertainty, as well as demonstrating how these barriers change as a function of voltage. We believe that our QCA methodology represents a higher accuracy method for determining these reaction barriers in realistic battery contexts as we control for the electronic voltage of the calculation.
3.2.2 Li transport in LiF-LiF grain boundary
Finally, we examine the reaction barrier for a Li ion moving through the channel of the 5 LiF-LiF grain boundary explored earlier. This reaction barrier additionally includes the interface with the Li anode, which can provide a key differentiating factor. This approach contrasts with previous work of Ramasubramanian et al. 10, which finds the barrier for Li diffusion through this grain boundary to be 0.68 eV when examining the SEI in isolation. Using standard NEB calculations where we include the Li anode interface, we predict a diffusion barrier of 0.17 eV, which is already significantly reduced. Once again, we then extend this analysis with QCA of the initial and climbing image to determine the diffusion barrier as a function of voltage. Using a linear regression, we predict the diffusion barrier magnitude as
Once again, the dependence on voltage appears fairly minimal, although the inclusion of electronic voltage control lowers the overall barrier by 0.07 eV.
Our work highlights that previous work that has not been including both interfaces with anodes and electronic voltage may significantly overestimate reaction barriers for Li ion transport. This overestimation seems largely driven by the lack of inclusion on an anode-SEI interface within the reaction barrier calculations. Altering the voltage on the material, however, can have impacts on the total energy needed for a Li to diffuse across the full SEI, and represents a useful tool for tuning a given devices performance. This demonstrates that some reaction barriers which had previously been theoretically predicted to be too high for realistic operation may actually be feasible when controlling for interface and voltage effects within the calculation.
@@ 4 Discussion
Our simulations are necessarily imperfect representations of real battery devices. While QCA extends the length scale examined perpendicular to the surface, it does not extend the lateral area explored, which is treated using typical DFT periodic boundary conditions. This means that simulations of individual SEIs such as LiF and Li2O inherently assume that only these perfectly crystalline chemistries exist in the SEI. In reality, the surface is known to likely be a mixture of different SEI chemistries interacting with each other. While our simulations of grain boundaries begin to address this concern, they are necessarily still periodic and are thus assuming a repeating pattern of grain boundaries across the surface, often with quite high surface densities. Our work will serve as a stepping stone, however, for future multiscale research aiming to address the larger lateral surface.
@@
Another aspect that limits the potential real world accuracy of our calculations is the representation of the electrolyte. For ease of demonstrating QCA within these systems, we only simulated rather shallow monolayers of DOL electrolyte molecules along with the inorganice SEI, but organic SEI components based on reactions with the electrolyte are often observed as well 42. This will be addressed with future work including a realistic mix of electrolyte molecules and a full Guoy-Chapman-Stern analytic inclusion of the charge distribution. QCA represents a tractable method for including as much information about electrolytes as feasible into these calculations before resorting to relatively expensive AIMD.
Finally, our work is predicated on the SEI behaving largely as an insulator. The input charge carrier concentration is assumed to be uniform throughout the system, but may also be viewed as an average concentration for these system. This does not, however, capture when there is significant local variation in this concentration, which determines whether a system is insulating or (semi)conducting. Future work will focus on integrating regions of varied conductivity into our calculations at the expense of additional DFT SEI calculations.
@@ 5 Conclusion
Our QCA analysis reveals that the barrier for transporting Li through the SEI of a battery is dependent on the overall applied electronic voltage, and to large extent on the location of the Li ion in the SEI. This implies that controlling the chemical composition of the SEI and applied voltage may be a useful tool for modulating the rate of charging for a given battery systems. In general, this trend matches intuition as higher open circuit voltages lead to lower transport barriers, favoring easier charging. The threshold voltage at which Li ion movement is encouraged toward the anode versus away from the anode, however, is highly dependent on the specific SEI chemistry. We see that, at least in some cases, grain boundaries lower this potential threshold more than would be seen for the individual components. We find that controlling for interface and electronic voltage effects can significantly lower the predicted Li diffusion barriers when compared to literature. This implies that some previous reactions which had been dismissed as requiring too much energy may be plausible when correctly accounting for these effects. Furthermore, it is likely that as models of the anode-SEI interface become more realistic, this effect will become more pronounced, as evidenced by the interface’s significant impact on our QCA calculations.
@@
Supporting information is available providing: (1) input files with the atomic structures of all systems calculated throughout this work, and (2) calculation of the dependence of the calculated work function on the number of explicit solvent layers included.